OpenStudy (theedwardsfamily):

Six friends together earn $167.88 selling crafts. How much money will each friend get if they share the money equally? Complete the equation to model the problem. Use m to represent the unknown value.

OpenStudy (jhannybean):

let m represent the friends. You have 6 of them, so 6m. They want to distribute the money equally, so \(6m = 167.88\)

OpenStudy (jhannybean):

Solving for m, you will find how much each friend gets.

OpenStudy (jhannybean):

Use what I have written to fill in the blanks.

OpenStudy (jhannybean):

Solve for m

OpenStudy (theedwardsfamily):

oook

OpenStudy (jhannybean):

Actually, don't solve it.

OpenStudy (jhannybean):

Just set up the fraction.

OpenStudy (jhannybean):

How will you isolate the m?

OpenStudy (jhannybean):

What do you have to do to the left side of the equation to get m by itself?

OpenStudy (theedwardsfamily):

um.....do I have to put 1676.88 in fraction?

OpenStudy (jhannybean):

That's part of it, but answer my question first.

OpenStudy (theedwardsfamily):

well this is tricky becase Im not a fan of math

OpenStudy (jhannybean):

ok, you have two options.

OpenStudy (jhannybean):

multiplying by 6 will only increase the value of the right side and will not get rid of the constant that is multiplying m

OpenStudy (jhannybean):

on the other hand you can divide out the 6 so you will get m by itself

OpenStudy (theedwardsfamily):

Still stuck

OpenStudy (theedwardsfamily):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (danjs):

\[\frac{ m ~dollars }{ 1 ~friend } = \frac{ 167.88~dollars }{ 6~friends }\]

OpenStudy (danjs):

m dollars is to 1 friend, as 167.88 dolalrs is to 6 friends

OpenStudy (danjs):

just write that in reverse order to put it like they have in those boxes, i just noticed that link

OpenStudy (danjs):

167.88 / 6 = m dollars per friend

OpenStudy (theedwardsfamily):

thanks! I understand:DDDDDDD

OpenStudy (danjs):

welcome